A firm uses skilled (S) and unskilled (U) labour to produce output (Q). The firm’s production function is given by

   "f(U,S)= sqrt(2U+4S)"

where U denotes unskilled and S denotes skilled labour. Assume U ≥ 0 and S ≥ 0. Let wu = 6 and ws = 8 denote the prices of unskilled labour and skilled labour respectively.

a. Suppose (U,S) = (8,5). How much output does the firm produce? What is the total cost of producing that level of output?

b. Let Q0 and C0 respectively denote the value of output and total costs obtained in part (a). Firm is not particularly interested in cost-cutting but they are averse to cost-increase. Find out the maximum level of output (Q*) that can be produced with a spending cap of  C0  and recommend the levels of unskilled (U∗) and skilled labour (S∗) necessary to produce Q∗. Find Q∗, U∗, and S∗.



1
Expert's answer
2021-03-16T12:01:50-0400

a.To maximize output at these costs, you can use the direct equal cost, or isocost

"C = P \times U+ P \times S"

Substituting the original data, we get the equation:

6U+8S=8.5

"U=\frac{8.5-8S}{6}"

Let's substitute this value into the given equation

"Q=\sqrt{2\times(\frac{8.5-8S}{6})+4S}=\sqrt{(\frac{8.5-8S}{3})+4S}=\sqrt{(\frac{8.5+4S}{3})}"

The aggregate product function reaches a maximum when the marginal product function is zero

"Q'=\frac{2\sqrt{6}}{3\sqrt{8S+17}}"

The marginal product is the first derivative of the aggregate product by a variable factor


 "\frac{2\sqrt{6}}{3\sqrt{8S+17}}=0"


the corresponding equation has no solution


b.6U+8S=0

"U=\frac{-8S}{6}=\frac{-4S}{3}"

"Q=\sqrt{2\times(\frac{-4S}{3})+4S}=\sqrt{(\frac{-8S}{3})+4S}=\sqrt{(\frac{-4S}{3})}"

"Q'=\frac{S}{\sqrt{3}\times (-S)^{1/5}}"

"\frac{S}{\sqrt{3}\times (-S)^{1/5}}=0"


S=0

U=0

Q=0